OpenStudy (anonymous):

Chords AB and CD intersect at E -If CE=12, ED=2, And AE=3, find EB

OpenStudy (anonymous):

EB- 11

OpenStudy (mertsj):

I believe the product of the segments should be equal.

OpenStudy (mertsj):

Just like bluepig said.

OpenStudy (anonymous):

THM: chords intersect in a circle -> product of parts equal.  3 * x = 2 * 12 x = 2 * 4 = 8 EB = 8